Find equivalent functional definition for each of the following and draw a graph of the function.


f(x) = |x|•|x-1|

1
Expert's answer
2022-06-12T15:41:50-0400
"x<0, f(x)=-x(1-x)=x^2-x"

"0\le x<1, f(x)=x(1-x)=-x^2+x"

"x\ge1, f(x)=x(x-1)=x^2-x"

"f(x)= \begin{cases}
 x^2-x &x<0 \\
 -x^2+x &0\le x<1\\
 x^2-x & x\ge1
\end{cases}"